For every 2 salt and vinegar crisps, there are 4 packets of plain crisps, 3 packets of cheese and onion crisps and 1 packet of prawn cocktail crisps.
How many salt and vinegar flavour packets are there if total number of packets is 50 ?
Kindly help :/

Answer by Theo(13342)

x = number of prawn
2x = number of salt
4x = number of plain
3x = number of cheese

x + 2x + 4x + 3x = 50

10x = 50

x = 5

number of prawn is 5.
number of salt is 10.
number of plain is 20.
number of cheese is 15.

add them up and you get 50.

problem statement says:

For every 2 salt and vinegar packets, there are 4 packets of plain crisps, 3 packets of cheese and onion crisps and 1 packet of prawn cocktail crisps.

for every 2 salt there are 4 plain.
4/2 = 2
there should be 2 times as many plain as salt.
there are because 20 is 2 times 10.

for every 2 salt there are 3 cheese.
3/2 = 1.5
there should be 1.5 times as many cheese as salt.
there are because 15 = 1.5 * 10.

for every 2 salt there is 1 prawn.
1/2 = .5
there should be .5 times as many prawn as salt.
there are because 5 is .5 * 10.

looks good.

you were asked how many salt and vinegar?
answer is 10.